# Open Vocabulary Semantic Scene Sketch Understanding

### v1

@@ Abstract
We study the underexplored but fundamental vision problem of machine understanding of abstract freehand scene sketches. We introduce a sketch encoder that results in semantically-aware feature space, which we evaluate by testing its performance on a semantic sketch segmentation task. To train our model we rely only on the availability of bitmap sketches with their brief captions and do not require any pixel-level annotations. To obtain generalization to a large set of sketches and categories, we build on a vision transformer encoder pretrained with the CLIP model. We freeze the text encoder and perform visual-prompt tuning of the visual encoder branch while introducing a set of critical modifications. Firstly, we augment the classical key-query (k-q) self-attention blocks with value-value (v-v) self-attention blocks. Central to our model is a two-level hierarchical network design that enables efficient semantic disentanglement: The first level ensures holistic scene sketch encoding, and the second level focuses on individual categories. We, then, in the second level of the hierarchy, introduce a cross-attention between textual and visual branches. Our method outperforms zero-shot CLIP pixel accuracy of segmentation results by 37 points, reaching an accuracy of on the FS-COCO sketch dataset. Finally, we conduct a user study that allows us to identify further improvements needed over our method to reconcile machine and human understanding of scene sketches.
@@ 1 Introduction
Freehand sketches can be represented as a sequence or cloud of individual strokes, or as a bitmap image. As one of the first works on scene sketch understanding, we target a general setting where we assume only the availability of bitmap representations.
We also aim at the method that can generalize to a large number of scenes and object categories. To this end, we build our sketch encoder on a Visual Transformer (ViT) encoder pre-trained with a popular CLIP [43] foundation model (Fig. 1). We propose a two-level hierarchy in our network, where the two levels (“Holistic” and “Category-level”) share the weights of our visual encoder. The first level focuses on ensuring that our model can capture holistic scene understanding (Fig. 2: I. Holistic), while the second level ensures that the encoder can efficiently encode and distinguish individual categories (Fig. 2: II. Category-level). We avoid reliance on tedious user per-pixel annotations by leveraging sketch-caption pairs from the FS-COCO dataset [8], and aligning the visual tokens of sketch patches with textual tokens from the sketch captions, using triplet loss training. We strengthen the alignment by introducing sketch-text cross-attention in the second level of the network’s hierarchy (Fig. 2: g.). Additionally, we introduce a modified self-attention computation to the visual transformer encoder used in both layers, inspired by recent work by Li et al. [32].
We conduct a comprehensive evaluation of our method comparing it with recent language-supervised image segmentation methods [58, 43, 32], fine-tuned on the FS-COCO dataset. We show that our approach outperforms with a large margin all existing methods on the task of freehand sketch segmentation. We also compare with a previous fully supervised work on scene sketch segmentation [19], trained on a semi-synthetic set of sketches composed of individual category sketches. We demonstrate that their work does not generalize well to freehand scene sketches from the FS-COCO dataset. Our method demonstrates consistent performance and similarly outperforms [19] on a dataset of freehand sketches provided by Ge et al. [19].
@@
In summary, our contributions include: (1) a two-level network design used during training, focusing on holistic scene sketch understanding and category disentanglement, (2) the first language-supervised scene sketch segmentation method, (3) per pixel segmentation annotations of 975 sketches from the FS-COCO dataset, and (4) multi-user annotations of a subset of distinct groups of sketches.
@@ 2.1 Unsupervised and weekly supervised image semantic segmentation
The need for pixel-wise segmentation limits the amount of data that fully supervised segmentation models [35, 64, 5, 6, 1, 17] can use for training, as such annotations are costly to collect. This in its turn limits the generalization properties of models trained with pixel-level annotations. To avoid the need for extensive annotations, unsupervised [62, 7, 23, 25, 38], semi-supervised [39, 72] and weakly supervised [40, 56, 58, 37, 13, 12, 36, 69] methods were proposed.
Our method belongs to the group of weakly supervised methods based on text annotations only [58, 37, 13, 12, 36, 4], such methods are not limited to a fixed set of categories and therefore are referred to as open vocabulary semantic segmentation methods. Image methods typically rely on the spatial proximity of semantically similar pixels. This is less applicable in the sparse and largely monochromatic landscape of freehand sketches. For example, recent GroupViT [58] and SegCLIP [37] use learnable group tokens and semantic group modules to aggregate low-layer pixel features. In our work, we propose a two-level training architecture taking sketch specifics into account.
@@ 2.2 Sketch semantic segmentation
Semantic scene sketch segmentation [51], and more broadly scene sketch understanding, is underexplored to a large extent due to lack of data. The lack of data is typically addressed by introducing semi-synthetic sketch datasets. The SketchyScene dataset [73] consists of 7,264 sketch-image pairs, obtained by arranging clip-art sketches in alignment with a reference image. SketchyCOCO dataset [18] is generated from COCO-Stuff [3] by semi-automatically arranging freehand sketches of individual categories. Ge et al. [19] introduced their own semi-synthetic scene sketch dataset and adopted a DeepLab-v2 [5] architecture to the scene sketch segmentation task. SketchSeger [59] proposed an encoder-decoder model based on hierarchical Transformers, trained with a stroke-based cross-entropy loss on semi-synthetic scene sketches formed by combining sketches from the QuickDraw dataset [21]. Zhang et al. [63] proposed an RNN-GCN-based architecture trained on annotated freehand scene sketches. However, neither the dataset nor the code have been released. We do not require stroke-level information or pixel-wise segmentation of the training data, and leverage the FS-COCO dataset [8] of freehand sketches with their textual descriptions.
@@ 3.1.1 Visual encoder
Attention computation in the visual encoder
It was observed by Li et al. [32] that CLIP-predicted similarity maps between image and text features emphasize background regions rather than areas that correspond to a category in the text embedding.
To address this issue, they proposed an alternative configuration of the visual transformer that does not require training or fine-tuning the original model. They demonstrated that this configuration leads to improved performance in open vocabulary segmentation tasks.
We performed a similar experiment with CLIP features for sketch inputs: The similarity maps in the second row of Fig. 3 show the poor ability of CLIP features to identify target categories.
Therefore, we follow Li et al. [32] and use their two-path configuration of the vision transformer. However, we use it not only for inference, but also incorporate this two-path configuration directly into our network training, as we find it more beneficial. We provide a detailed analysis in Sec. 4.5.1.
@@ 3.2 Categories disentanglement
Given the sketch caption we automatically identify individual categories and generate a set of textual prompts of the form “A sketch of *” (Fig. 2b.). Each of these textual category prompts is encoded with the CLIP text encoder into (Caption Category Token).
@@
where is the patch index and is an index of a category (e.g. tree). The resulting similarity matrix represents the category label probabilities for each individual patch (Fig. 2d.). To generate a pixel-level similarity map, we reshape each and then upscale to the dimensions of the original scene sketch using bi-cubic interpolation [52]. By multiplying these per category maps with the input scene sketch, as shown in Fig. 2e., we obtain disentanglement into individual sketch categories.
3.2.1 Thresholding with a learnable parameter
@@
3.2.2 Visual Encoder with Cross Attention
@@
3.2.3 Text-sketch category-level alignment
@@ 3.4 Inference
If we want to isolate just a few categories in the sketch, we can use the thresholding strategy that we use during training to isolate the pixels of a given category (Fig. 2f.). We used this strategy to obtain visualizations in Fig. 1, with a threshold value of that we found to be optimal on the test set of sketches. We do not use the learned value from the training, as during training the model does not have to select all the pixels of the given category, but only those that are sufficient to confidently predict the category label. We provide in-depth discussion in the supplemental.
@@ 4.1 Training and test data
Test/Validation set annotations One of the co-authors manually annotated test and validation sketches, relying on reference images and category labels from the MS-COCO [33] dataset. We assign each stroke a unique category label. Candidate category labels are extracted from MS-COCO image captions rather than sketch captions to obtain richer ‘ground-truth’ annotations. Our test set contains 185 different object classes, with an average of 3.54 objects per sketch.
@@ 4.4 Comparison against state-of-the-art
4.4.1 Comparison with methods relying on the availability of pixel-level annotations
We first compare with several recent methods for image segmentation that similarly to us utilize either CLIP as a backbone: DenseCLIP [44] and ZegCLIP [69], or more recent foundational backbones Grounding-DINO [34] and SAM [28], used in Grounded-SAM [20]. These methods require a certain amount of pixel-level annotated examples. We also compare to a recent fully supervised method [19] for scene sketch semantic segmentation, which is trained on a dataset of semi-synthetic sketches. Such sketches are obtained as a superposition of freehand category-level sketches. We refer to this method as LDP (Local Detail Perception).
Tab. 1 shows that neither of the considered methods generalizes well to freehand scene sketches with a large number of categories.
@@ 4.4.2 Comparison with language-supervised methods
Table 2 shows that our method outperforms all considered baselines, and surpasses the best-performing baseline CLIP Surgery by a substantial margin of , and points in score, and , respectively. In Sec. 4.5.1, we evaluate various elements of our architecture and their contribution to overall performance.
| Methods | |||
| CLIP [43] | |||
| GroupViT [58] | |||
| SegCLIP [37] | |||
| CLIP_Surgery [32] | |||
| CLIP | |||
| GroupViT | |||
| SegCLIP | |||
| CLIP_Surgery | |||
| CLIP_Surgery | |||
@@ 4.4.3 Generalization ability of our method
Here, we evaluate our method on an additional dataset of 50 freehand sketches provided and annotated by Ge et al. [19]. Tab. 3 shows that our model again demonstrates superior performance on this dataset over the method [19], fully supervised on semi-synthetic sketches. We do not compute Acc@S as sketches are only available as bitmap images. This experiment highlights that short language captions can be efficiently used for training, eliminating the need for expensive and time-consuming per-pixel annotations.
The lower mIoU values on these sketches than on FS-COCO sketches can be explained by (1) on larger average number of categories in them ( categories per sketch) than in our FS-COCO test set ( categories per sketch); (2) domain gap. The sketches from [19] contain symbolic representations of objects (see inset
@@ 5.2 User study setting
Participants We recruited 25 participants ( male) for this study. Each participant was randomly assigned six sketches: 1 from each of the first 4 groups and 2 from the 5th group, such that every sketch was annotated by five unique participants.
@@ 5.3 User study analysis and Future work
Interestingly, given a label humans can correctly identify sketch pixels even in the presence of ambiguity (Fig. 8 Ambiguous). While none of the models currently match human performance on ambiguous sketches, our model surpasses the other methods by a noticeable margin, demonstrating the effectiveness of our two-level training architecture.
@@ 6 Conclusion
S2 Additional analysis of our method performance
@@
In this section, we analyze segmentation accuracy per category in both the train and test sets. We show in Fig. S11 the pixel accuracy (Acc@P) for each selected object category. For the figure, we selected categories that appear more than ten times in FS-COCO dataset [8] captions. First, we can see that the segmentation accuracy is smoothly distributed across different categories.
Next, we investigate whether more frequent categories are more likely to be labeled accurately. To evaluate this, we approximate the frequency of a category by counting its occurrence in both the train and test sets, then consider only categories that appear in the test set. We plot with a green and red lines in Fig. S11 the train and test sets category frequency, respectively.
@@
S2.3 Additional discussion on synthetic vs. freehand sketches
@@
Fine-tuning on semi-synthetic sketches
We perform two additional experiments:
1.
2.
The results are shown in Tab. S5: Ours and Ours.
@@
In the main paper, in Sec. 5.1, we introduced four challenging categories of sketches for our method that we used for the user study. We show all the sketches used in the user study in Fig. S12. For convenience, below we repeat the definition of each category:
@@
S3.2 Visual analysis of the segmentation results for the sketches with “Interchangeable categories” obtained with our model and from the user study
@@
As described in Sec. 4.3 of the main paper, for each of the models fine-tuned on sketch data: ours and competing methods, we select a checkpoint based on the performance on the validation set with pixel-lelve segmentation annotations, consisting of 475 sketches. This requires at training time having a small set of pixel-level annotated sketches, which can be limiting. However, we observe that the loss gradually decreases for our model, and it is safe to choose a last checkpoint if such an annotated set is not available. In Tab. S9, we provide a comparison with the results when for our model and competing models the last checkpoint is used. We trained for 20 epochs. We observe that after that the convergence rate is very low for each of the considered models.
@@
We also observe that with either of the choices of a checkpoint, the performance on the validation and test sets is similar, with just a small decrease in performance on the test set compared to the validation set. Our test set includes sketches from five non-expert artists whose sketches were not present in either the training or validation sets. Therefore, this analysis implies that there is no over-fitting to the training data and our model robustly generalizes to the unseen skecthes and drawing styles.

### v2

@@ Abstract
We study the underexplored but fundamental problem of machine understanding of abstract freehand scene sketches. We introduce a sketch encoder that ensures a semantically-aware feature space, which we evaluate by testing its performance on a semantic sketch segmentation task. To train our model, we rely only on bitmap sketches accompanied by brief captions, avoiding the need for pixel-level annotations. To generalize to a large set of sketches and categories, we build upon a vision transformer encoder pretrained with the CLIP model. We freeze the text encoder and perform visual-prompt tuning of the visual encoder branch while introducing a set of critical modifications. First, we augment the classical key-query (k-q) self-attention blocks with value-value (v-v) self-attention blocks. Central to our model is a two-level hierarchical training that enables efficient semantic disentanglement: The first level ensures holistic scene sketch encoding, and the second level focuses on individual categories. In the second level of the hierarchy, we introduce cross-attention between the text and vision branches. Our method outperforms zero-shot CLIP segmentation results by 37 points, reaching a pixel accuracy of on the FS-COCO sketch dataset. Finally, we conduct a user study that allows us to identify further improvements needed over our method to reconcile machine and human understanding of freehand scene sketches.
@@ 1 Introduction
Freehand sketches can be represented as a sequence or cloud of individual strokes, or as a bitmap image. As one of the first works on scene sketch understanding, we target a general setting where we assume only the availability of bitmap representations. We also aim at the method that can generalize to a large number of scenes and object categories. To this end, we build our sketch encoder on a Visual Transformer (ViT) encoder pre-trained with a popular CLIP [46] foundation model (Fig. 1). We propose a two-level hierarchical training of our network, where the two levels (“Holistic" and “Category-level") share the weights of our visual encoder. The first level focuses on ensuring that our model can capture holistic scene understanding (Fig. 2: I. Holistic), while the second level ensures that the encoder can efficiently encode and distinguish individual categories (Fig. 2: II. Category-level). We avoid reliance on tedious user per-pixel annotations by leveraging sketch-caption pairs from the FS-COCO dataset [10], and aligning the visual tokens of sketch patches with textual tokens from the sketch captions, using triplet loss training. We strengthen the alignment by introducing sketch-text cross-attention in the second level of the network’s hierarchy (Fig. 2: g.). Additionally, we introduce a modified self-attention computation to the visual transformer encoder used in both layers, inspired by recent work by Li et al. [35].
We conduct a comprehensive evaluation of our method comparing it with recent language-supervised image segmentation methods [61, 46, 35], fine-tuned on the FS-COCO dataset. We show that our approach outperforms with a large margin all existing methods on the task of freehand sketch segmentation. We also compare with a previous fully supervised work on scene sketch segmentation [21], trained on a semi-synthetic set of sketches composed of individual category sketches. We demonstrate that their work does not generalize well to freehand scene sketches [10]. Our method demonstrates consistent performance and similarly outperforms [21] on a dataset of freehand sketches provided by Ge et al. [21].
@@
In summary, our contributions include: (1) a two-level hierarchical training approach, focusing on holistic scene sketch understanding and category disentanglement, (2) the first language-supervised scene sketch segmentation method, (3) per pixel segmentation annotations of 975 sketches from the FS-COCO dataset, and (4) multi-user annotations of a subset of distinct groups of sketches.
@@ 2.1 Unsupervised and Weekly Supervised Image Semantic Segmentation
The need for pixel-wise segmentation limits the number of instances that supervised segmentation models [38, 67, 7, 8, 1, 19] can use for training, as such annotations are costly to collect. This in turn limits the generalization properties of models trained with pixel-level annotations. To avoid the need for extensive annotations, unsupervised [65, 9, 25, 28, 41], semi-supervised [42, 75] and weakly supervised [43, 59, 61, 40, 15, 14, 39, 72, 26] methods were proposed.
Our method belongs to the group of weakly supervised methods based on text annotations only [61, 40, 15, 14, 39, 6], such methods are not limited to a fixed set of categories and therefore are referred to as open vocabulary semantic segmentation methods. Image methods typically rely on the spatial proximity of semantically similar pixels. This is less applicable in the sparse and largely monochromatic landscape of freehand sketches. For example, recent GroupViT [61] and SegCLIP [40] use learnable group tokens and semantic group modules to aggregate low-layer pixel features. In our work, we propose a two-level training architecture taking sketch sparsity and abstraction into account.
@@ 2.2 Sketch Semantic Segmentation
Semantic scene sketch segmentation [54], and more broadly scene sketch understanding, is underexplored, to a large extent due to a lack of data. The lack of data is typically addressed by introducing semi-synthetic sketch datasets. The SketchyScene dataset [76] consists of 7,264 sketch-image pairs, obtained by arranging clip-art individual category sketches in alignment with a reference image. SketchyCOCO dataset [20] is generated from COCO-Stuff [4] by semi-automatically arranging freehand sketches of individual categories. Ge et al. [21] introduced their own semi-synthetic scene sketch dataset and adopted a DeepLab-v2 [7] architecture to the scene sketch segmentation task. SketchSeger [62] proposed an encoder-decoder model based on hierarchical Transformers, trained with a stroke-based cross-entropy loss on semi-synthetic scene sketches formed by combining sketches from the QuickDraw dataset [23]. Zhang et al. [66] proposed an RNN-GCN-based architecture trained on annotated freehand scene sketches. However, neither the dataset nor the code have been released. We do not require stroke-level information or pixel-wise segmentation of the training data, and leverage the FS-COCO dataset [10] of freehand sketches with their textual descriptions.
@@ 3.1.1 Visual encoder
Attention computation
It was observed by Li et al. [35] that CLIP-predicted similarity maps between image and text features emphasize background regions rather than areas that correspond to a category in the text embedding. To address this issue, they proposed to use an instance of self-self attention called v-v attention, which does not require training or fine-tuning the original model. Li et al. [35], and later Bousselham et al. [3] demonstrated that this leads to improved performance in open vocabulary segmentation tasks: Self-self-attention reinforces the similarity of tokens already close to each other (e.g. representing the same object), which leads to a clearer separation in the feature space, thereby improving the segmentation quality.
We performed a similar experiment with CLIP features for sketch inputs: The similarity maps in the second row of Fig. 3 show the poor ability of CLIP features to identify target categories. Therefore, we follow [35] and use their two-path configuration of the vision transformer. However, we use it not only for inference but also incorporate this two-path configuration directly into our network training, as we find it more beneficial. We provide a detailed analysis in Sec. 4.5.1.
@@ 3.2 Categories Disentanglement
Given the sketch caption we automatically identify individual categories and generate a set of textual prompts of the form “A sketch of *" (Fig. 2b.). Each of these textual category prompts is encoded with the CLIP text encoder into (Caption Category Token).
@@
where is the patch index and is an index of a category (e.g. trees). The resulting similarity matrix represents the category label probabilities for each individual patch (Fig. 2d.). To generate a pixel-level similarity map, we reshape each and then upscale to the dimensions of the original scene sketch using bi-cubic interpolation [55]. By multiplying these per category maps with the input scene sketch, as shown in Fig. 2e., we obtain disentanglement into individual sketch categories.
Thresholding with a learnable parameter
@@
Visual encoder with Cross-Attention
@@
Text-sketch category-level alignment
@@ 3.4 Inference
If we want to isolate just a few categories in the sketch, we can use the thresholding strategy that we use during training to isolate the pixels of a given category (Fig. 2f.). We used this strategy to obtain visualizations in Fig. 1, with a threshold value of that we found to be optimal on the test set of sketches. We do not use the learned value from the training, as during training the model does not have to select all the pixels of the given category, but only those that are sufficient to confidently predict the category label. We provide an in-depth discussion in the supplemental.
@@ 4.1 Training and Test Data
Annotations One of the co-authors manually annotated test and validation sketches, relying on reference images and category labels from the MS-COCO [36] dataset. We assign each stroke a unique category label. Candidate category labels are extracted from MS-COCO image captions rather than sketch captions to obtain richer ‘ground-truth’ annotations. Our test set contains 185 different object classes, with an average of 3.54 objects per sketch.
@@ 4.4 Comparison against state-of-the-art
4.4.1 Comparison with fully-supervised methods
We first compare with several recent methods for image segmentation that similarly to us utilize either CLIP as a backbone: DenseCLIP [47] and ZegCLIP [72], or more recent foundational backbones Grounding-DINO [37] and SAM [31], used in Grounded-SAM [22]. These methods require pixel-level annotated examples, and therefore can not be fine-tuned on our training data. We also compare to a recent fully supervised method LDP (Local Detail Perception) [21] for scene sketch semantic segmentation, which is trained on a dataset of semi-synthetic sketches. Such sketches are obtained as a superposition of freehand category-level sketches. Tab. 1 shows that neither of the these methods generalizes well to freehand scene sketches.
@@ 4.4.2 Comparison with language-supervised methods
Tab. 2 shows that our method outperforms all considered baselines, and surpasses the best-performing baseline CLIP Surgery by a substantial margin of , and points in score, and , respectively. In Sec. 4.5.1, we evaluate various elements of our architecture and their contribution to overall performance.
| Methods | ||||
| Zero-shot | CLIP [46] | |||
| GroupViT [61] | ||||
| SegCLIP [40] | ||||
| CLIP_Surgery [35] | ||||
| Fine-tuned | CLIP | |||
| GroupViT | ||||
| SegCLIP | ||||
| CLIP_Surgery | ||||
| CLIP_Surgery | ||||
@@ 4.4.3 Generalization ability of our method
Next, we evaluate our method on an additional dataset of 50 freehand sketches provided and annotated by Ge et al. [21]. Tab. 3 shows that our model again demonstrates superior performance on this dataset over the method [21], fully supervised on semi-synthetic sketches. We do not compute Acc@S as sketches are only available as bitmap images. This experiment highlights that short language captions can be efficiently used for training, eliminating the need for expensive and time-consuming per-pixel annotations.
The lower mIoU values on these sketches than on FS-COCO sketches can be explained by (1) on larger average number of categories in them ( categories per sketch) than in our FS-COCO test set ( categories per sketch); (2) domain gap. The sketches from [21] contain symbolic representations of objects (see the inset
@@ 5.2 User Study Setting
Participants We recruited 25 participants ( male). Each participant was randomly assigned 6 sketches: 1 from each of the first 4 groups and 2 from the 5th group, such that every sketch was annotated by five unique participants.
@@ 5.3 User Study Analysis and Future Work
Fig. 8 Ambiguous shows that, given a label, humans can correctly identify sketch pixels even in the presence of ambiguity. While none of the models currently match human performance on ambiguous sketches, our model surpasses the other methods by a noticeable margin, demonstrating the effectiveness of our two-level training architecture.
@@ 6 Conclusion
In Sec. S5, we provide the computational cost of our method.
S2 Additional Performance Analysis
@@
In this section, we analyze segmentation accuracy per category in both the train and test sets. We show in Fig. S11 the pixel accuracy (Acc@P) for each selected object category. For the figure, we selected categories that appear more than ten times in the FS-COCO dataset [10] captions. First, we can see that the segmentation accuracy is smoothly distributed across different categories.
Next, we investigate whether more frequent categories are more likely to be labeled accurately. To evaluate this, we approximate the frequency of a category by counting its occurrence in both the train and test sets, then consider only categories that appear in the test set. We plot with green and red lines in Fig. S11 the train and test sets category frequency, respectively.
@@
Model generalization to new object categories
Our test set includes 185 object classes, with 125 seen and 60 unseen during training. The accuracy on seen categories is and on unseen. These results demonstrate good generalization of our model to unseen categories.
S2.3 Synthetic vs. Freehand sketches
@@
| Method | |||
| LDP [21] | |||
| Ours | |||
| Ours | |||
| Ours |
Fine-tuning on semi-synthetic sketches
We perform two additional experiments:
1.
2.
The results are shown in Tab. S5: Ours and Ours.
@@
S2.3.2 Pre-training on synthetic sketches
We also experiment with fine-tuning CLIP and CLIPSurgery on synthetic sketches. However, training on millions of synthetic sketches is out of the scope of this work due to computational constraints. As a feasible experiment, we generated synthetic sketches for the reference images in our training set, using [5], in ‘contour’ style (as the closest to the test set sketches style). This is the number of sketches identical to the number of sketches we use to train our model. The accuracy on our test set of fine-tuned this way CLIP and CLIPSurgery increases by negligible to points compared to their zero-shot performance. In comparison, our model outperforms their zero-shot performance by and points, respectively. Training our model from CLIP weights pre-trained on synthetic sketches boosts the performance only by points.
@@
In the main paper, in Sec. 5.1, we introduced four challenging categories of sketches for our method, that we used for the user study. We show all the sketches used in the user study in Fig. S12. For convenience, below we repeat the definition of each category:
@@
S3.2 Annotators
We recruited 25 participants ( male). The annotators are PhD students in diverse disciplines and of diverse nationalities aged from 22 to 42 years (average age 29.32). We believe this group represents well the general population and each individual performed the task carefully.
S3.3 Visual Analysis of Interchangeable Categories Segmentation Results
@@
S3.4 Statistical Significance: Ours vs CLIPSurgery
On the 20 sketches from the 4 challenging groups, our model outperforms CLIPSurgery with a p-value of . In the ‘strong’ group, we have 10 sketches, in which, while our model performs on par with humans, it outperforms CLIPSurgery with a p-value of .
@@
As described in Sec. 4.3 of the main paper, for each of the models fine-tuned on sketch data: ours and competing methods, we select a checkpoint based on the performance on the validation set with pixel-level segmentation annotations, consisting of 475 sketches. This requires at training time having a small set of pixel-level annotated sketches, which can be limiting. However, we observe that the loss gradually decreases for our model, and it is safe to choose a last checkpoint if such an annotated set is not available. In Tab. S9, we provide a comparison with the results when for our model and competing models the last checkpoint is used. We trained for 20 epochs. We observe that after that the convergence rate is very low for each of the considered models.
@@
We also observe that with either of the choices of a checkpoint, the performance on the validation and test sets is similar, with just a small decrease in performance on the test set compared to the validation set. Our test set includes sketches from five non-expert artists whose sketches were not present in either the training or validation sets. Therefore, this analysis implies that there is no over-fitting to the training data and our model robustly generalizes to the unseen sketches and drawing styles.
| Training | Parameters | Inference | |||||
| GFLOPS | Hours | Epochs | # Trainable | # Full | GFLOPS | Secs. | |
| CLIP | 189.47 | 3.76 | 20 | 149M | 149M | 89.02 | 0.21 |
| CLIPSurgery | 231.41 | 4.08 | 20 | 162M | 162M | 113.64 | 0.73 |
| Ours | 346.36 | 7.31 | 20 | 16M | 165M | 121.59 | 1.05 |
@@
S5 Computational Cost
We detail in Tab. S10 the computational cost of our method compared to CLIP and CLIPSurgery. Our two-level hierarchical network design introduces additional complexity, through value-value self-attention and cross-attention blocks. However, we maintain a comparable level of complexity to CLIPSurgery during inference. This slight computational increase is justified given the substantial 13 mIoU points improvement over CLIP_Surgery (as shown in Tab. 4 in the main paper). Our code can be further optimized to reach the performance of CLIPSurgery at inference time.